[21] 2. Suppose Michigan Angels Group has already invested $3M in MotorCity Co. in return for 1M shares, resulting in the founders owning 70% of their company. In 2 years, additional funding is needed in the amount of $1M which is supplied by Mobility Ventures and this round is priced at $2 per share. Assume no ratchet provision.

a)  What is the number of shares of founders after the angel round?

b)  What is the total number of outstanding shares in MotorCity (owned by founders and angels)?

c)  What is proportion of shares owned by Mobility Ventures after the new (second) round?

d)  What is the proportion of shares owned by Michigan Angels after the new round?

e) What is proportion of shares owned by founders after the new round?

Assuming a full ratchet provision is in the term sheet, answer d and e.

g)  What is the ownership of Michigan Angels?

h) What is ownership of Mobility Ventures?
Answer #1

The Michigan Angels Group (MAG) has invested $3M in MotorCity Co. (MCC) in return of 1M shares. This implies that MAG acquired 1M shares at a price of $2 per shares ($3M/ 1M).

It is given that after the transaction the founder owns 70% stake in MCC, implying that the investor, MAG is owning 30% stake, i.e. 100% minus founders' stake.

Answer to (a).

No. of shares held by founders after MAG round = (Shares held by MAG/ Stake held by MAG) * Stake held by Founders = (1M/ 30%)* 70%  = 2.33M shares

Answer to (b)

No. of outstanding shares in MCC = No. of shares held by Founders + No. of shared held by MAG = 1M + 2.33M = 3.33M

Answer to (c)

To find = Proportion of shares owned by Mobility Ventures after the new (second) round.

No. of shares issued to Mobility Ventures (MV) = Funding supplied by MV/ Price per Share = $1M/ $2 per share = 0.5M shares

Total number of shares outstanding in MCC after the new round = total no. of shares before the new round + no. of shares issued to MV = 3.33M + 0.5M = 3.83M

Therefore, the proportion of shares owned by MV after the new (second) round = no. of shares held by MV/ total no. of shares outstanding = 0.5M/ 3.83M = 13.05%

Answer to (d)

Proportion of shares owned by Michigan Angels after the new round = No. of shares held by MAG/ total no. of outstanding shares = 1M/ 3.83M = 26.12%

Answer to (e)

Proportion of shares owned by Founders after the new round = No. of shares held by Founders/ total no. of outstanding shares = 2.33M/ 3.83M = 60.83%

Answer to (f)

The full ratchet provision in the term sheet protects the interest of earlier investor from anti-dilution of their stakes in case the new investor buys stake in the company at a lower valuation (price per share) as against the earlier investor.

MAG's purchase price per share = Purchase price/ No. of shares issued to MAG = $3M/ 1M = $3 per share

MV's purchase price share = $2 per share (given in the question)

The inclusion of full ratchet provision in the term sheet will protect MAG's interest and the shares allotted to MAG will get re-adjusted at the pricing given to the new investor, i.e MV

The adjusted shares owned by MAG, Founder and MV will be as follows,  

Shares held by MAG = Purchase price/ $2 per share =   $3M /$2 = 1.5M shares

Shares held by Founder = 2.33M (No change)

Shares held by MV = 0.5M (calculated earlier)

Total no. shares outstanding at MCC = 1.5M + 2.33M + 0.5M = 4.33M

Therefore,

Proportion of shares owned by MAG after the new (second) round = 1.5M/ 4.33M = 34.64%

Proportion of shares owned by Founders after the new (second) round = 2.33M/ 4.33M = 53.81%

Answer to (g)

Ownership of MAG = 1.5M/ 4.33M = 34.64%   [explained earlier]

Answer to (h)

Ownership of MV = Shares held by MV/ Total shares outstanding = 0.5M/ 4.33M = 11.55%